# Multi-marginal temporal Schrödinger Bridge Matching for video generation from unpaired data

### v1

@@ B.3 Critical Implementation Considerations
B.3.1 Scalability with High Dimensions and Many Marginals
@@
B.3.2 Masking and Time Discretization
@@
B.3.3 Interpolation Operator and Losses
@@
B.3.4 Time-Dependent Drift Networks
@@
B.4 Experiments details
The Adam optimizer is used with a learning rate of $2*10^{-4}$, and SiLU activations are applied on each layers unless stated otherwise. All experiments are executed on computing clusters that utilize GPU resources.
B.4.1 8Gaussians and Moons experiment
We used the same experimental setting as (Shi et al., 2023), except that we increase the batch size proportionally to the number of intermediate bridges. The 2-Wasserstein distance are computed with pot and the integrated path energy are computed with $\mathbb{E}\left[\int_{0}^{T}\|v(t,\mathbf{Z}_{t})\|^{2}\,dt\right]$ where $Z_{t}$ is the process simulated along the ODE drift 10.
@@ B.4 Experiments details
B.4.3 100d transcriptomic experiments
The dataset comprises 5 timepoints (Day 0 to Day 24) covering the progression from a homogeneous stem-cell population toward mesoderm, endoderm, and ectoderm precursors. The embryoid body dataset thus constitutes a realistic and challenging testbed for Schrödinger bridge methods, combining high dimensionality, non-Gaussian distributions, and branching lineages. We preprocessed the data following (Tong et al., 2020).
@@
We trained a network of about $3\times 10^{5}$ parameters for 20 outer iterations with 20,000 inner iterations.
B.4.4 MNIST experiments
@@
B.4.5 Biotine Experiments
@@
B.5 Proofs
Proof of variationnal proposition in Definition 3.2 (variational characterization).
By the additive property of the KL divergence (Léonard, 2014), for any $P\in\mathcal{P}(C([0,T],\mathbb{R}^{d}))$ and $\Pi\in\mathcal{R}^{\otimes}(\mathbb{Q})$, we can write
@@ B.5 Proofs
is convex and closed under the weak topology. Since the functional $P\mapsto D_{\mathrm{KL}}(P\|Q)$ is strictly convex, there is at most one minimizer.
@@
Therefore, (MMSB) admits a unique solution $P^{\star}$.
@@
The argument is identical to Proposition 2.10 in Léonard (2014), extended to the multi-marginal setting. For any admissible path measure $P\ll Q$, the additivity property of the relative entropy gives
@@
We follow the argument of (Léonard, 2014, Prop. 2.10). Fix an intermediate time $t_{k}$ with $0<k<n$. For any $Q\in\mathcal{P}(\Omega)$ and $z\in X$, set
@@
Let $\mu\in\mathcal{P}(X)$ and for each $z\in X$ prescribe $Q^{<}_{z}\in\mathcal{P}(\Omega_{[0,t_{k}]}\cap\{X_{t_{k}}=z\})$, $Q^{>}_{z}\in\mathcal{P}(\Omega_{[t_{k},1]}\cap\{X_{t_{k}}=z\})$. By the entropy additivity property (see formula (A.8) in Léonard (2014)), the measure
@@
The argument is a direct extension of Theorem 2.8 and Proposition 2.10 in Léonard (2014).
@@
and that they are internal in the sense of Proposition 2.6 of (Léonard, 2014).
@@
For the Markovian part, the equality follows analogously to the proof of (Shi et al., 2023).
@@
By integrating w.r.t. $\mathbb{P}$ and applying Csiszár’s Pythagorean identity (Csiszár, 1975, Eq. 2.6), we obtain
@@
Proof of Proposition 3.8.
As a reminder, we follow the same argument as in (Shi et al., 2023) and (De Bortoli et al., 2021). Applying Lemma 3.1, for any $N\in\mathbb{N}$ we obtain
@@
and boundedness $KL(\mathbb{P}^{n}\,\|\,\mathbb{P}^{\star})\leq KL(\mathbb{P}^{0}\,\|\,\mathbb{P}^{\star})<\infty.$ This proves the claim. ∎
Proof of Theorem 3.2.
As a reminder, the argument is the same as in (Shi et al., 2023) and (De Bortoli et al., 2021), but adapted to the multi-marginal setting.
By lower semicontinuity of KL, this implies
$\lim_{n\to\infty}KL(\mathbb{P}^{n}\,\|\,\mathbb{P}^{\star})=0.$
Finally, the inequality
Proof of Proposition 3.8.
We proceed by induction, adapting the argument of (De Bortoli et al., 2021, Appendix C.8).
@@
and $P^{\star}$ is indeed the weak solution produced by the IMFF algorithm. ∎

### v2

@@ A.4 Critical Implementation Considerations
A.4.1 Scalability with High Dimensions and Many Marginals
@@
A.4.2 Masking and Time Discretization
@@
A.4.3 Interpolation Operator and Losses
@@
A.4.4 Time-Dependent Drift Networks
@@
A.5 Experiments details
The Adam (kingma2017adammethodstochasticoptimization) or AdamW (loshchilov2019decoupledweightdecayregularization) optimizer is used with a learning rate of $2*10^{-4}$, and SiLU activations are applied on each layers unless stated otherwise.
A.5.1 Exact OT between Gaussian mixtures
In Figure 6 we can see the (exact) ”glued” OT plan empirically computed with POT. Observe how the global trajectory transports each Gaussian component of the mixture to a single other Gaussian component of the next marginal, yielding paths without any crossing. Note that the true multi-marginal transport plan remains unknown even in this simple Gaussian mixture setting.
Figure 6: Here we computed the OT plan between each pair of adjacent marginals empirically, in red and black lines. This plan can serve as a good proxy for the true multi-marginal plan.
A.5.2 8Gaussians and Moons experiment
We used the same experimental setting as (shi2023diffusionschrodingerbridgematching), except that we increase the batch size proportionally to the number of intermediate bridges. The 2-Wasserstein distance are computed with pot and the integrated path energy are computed with $\mathbb{E}\left[\int_{0}^{T}\|v(t,\mathbf{Z}_{t})\|^{2}\,dt\right]$ where $Z_{t}$ is the process simulated along the ODE drift 10.
@@ A.5.1 Exact OT between Gaussian mixtures
A.5.4 100d transcriptomic experiments: Embryoid Body
The dataset comprises 5 timepoints (Day 0 to Day 24) covering the progression from a homogeneous stem-cell population toward mesoderm, endoderm, and ectoderm precursors. The Embryoid Body dataset thus constitutes a realistic and challenging testbed for Schrödinger bridge methods, combining high dimensionality, non-Gaussian distributions, and branching lineages. We preprocessed the data following (tong2020trajectorynetdynamicoptimaltransport).
@@
We trained a network of about $300k$ parameters for 20 outer iterations with 20,000 inner iterations.
We show in Table 5 the performance advantage of our method compared to an iterative algorithm such as chen2023deepmomentummultimarginalschrodinger.
DSBM (chen2023deepmomentummultimarginalschrodinger) MMtSBM (ours)
number of marginals 4 5 4 5
Train time 33min 44 min 20 min 32 min
Sampling time 2.00s 2.02s 2.00 s 2.00 s
Table 5: Training and sampling times for chen2023deepmomentummultimarginalschrodinger and MMtSBM (ours) in dimension 100.
A.5.5 100d transcriptomic experiments: MULTI
We reused the preprocessed data from tong2024simulationfreeschrodingerbridgesscore. We do not whiten it. We conducted a minimal sweep to select the best $\sigma$ ($0.3$). The network is a simple $3$-layers MLP with around $500k$ parameters and we employ $150$ discretization time steps in total. Metrics are computed over $1k$ true test samples vs $1k$ generated samples, where these generations themselves come from the previous test marginal. We trained $3$ models with different seeds for each left-out time (either $t=1$ or $t=2$, corresponding to days $3$ and $4$). Our reported standard deviation is the pooled variance of the best same-hyperparameters $\sigma=0.3$ models over $2$ groups, each group corresponding to a left-out time. Other papers seem to have reported the overall variance, which we think makes less sense given the structure of the problem.
Group Number of runs Mean Std
Leave-out & test $t=1$ 3 37.026 0.822
Leave-out & test $t=2$ 3 52.059 0.367
Global 6 44.542 0.637
Table 6: Per-group statistics with pooled standard deviation $s_{\text{pooled}}=\sqrt{\sum(n_{i}-1)s_{i}^{2}\;/\;\sum(n_{i}-1)}$, where $n_{i}$ and $s_{i}$ are the sample size and standard deviation of each group.
Group
[…]
A.5.6 MNIST digit morphing experiment
@@
A.5.7 Biotine cell culture experiment
@@
Dataset Biotine
Dimension $128\times 128\times 3=49,152$
Number of marginals 6
Training time 5 h
Number of epochs 5
Sampling time 32 s
Generated frames 602
Table 7: Training and sampling statistics for video generation on the Biotine dataset.
The model trains in only 5 hours and subsequently generates an entire 602-frame trajectory in just 32 seconds, demonstrating both low training cost and highly efficient sampling.
A.6 Proofs
A.6.1 Definition 3.2
Proof of variational proposition in Definition 3.2 (variational characterization).
By the additive property of the KL divergence (leonard2014survey), for any $P\in\mathcal{P}(C([0,T],\mathbb{R}^{d}))$ and $\Pi\in\mathcal{R}^{\otimes}(\mathbb{Q})$, we can write
@@ A.6 Proofs
is convex. Since the functional $P\mapsto D_{\mathrm{KL}}(P\|Q)$ is strictly convex, there is at most one minimizer.
@@
Therefore, (MMSB) admits a unique solution $P^{\star}$. ∎
A.6.4 Proposition 3.2
@@
The argument is identical to Proposition 2.10 in leonard2014survey, extended to the multi-marginal setting. For any admissible path measure $P\ll Q$, the additivity property of the relative entropy gives
@@
We follow the argument of (leonard2014survey, Prop. 2.10). Fix an intermediate time $t_{k}$ with $0<k<n$. For any $Q\in\mathcal{P}(\Omega)$ and $z\in X$, set
@@
Let $\mu\in\mathcal{P}(X)$ and for each $z\in X$ prescribe $Q^{<}_{z}\in\mathcal{P}(\Omega_{[0,t_{k}]}\cap\{X_{t_{k}}=z\})$, $Q^{>}_{z}\in\mathcal{P}(\Omega_{[t_{k},1]}\cap\{X_{t_{k}}=z\})$. By the entropy additivity property (see formula (A.8) in leonard2014survey), the measure
@@
The argument is a direct extension of Theorem 2.8 and Proposition 2.10 in leonard2014survey.
@@
and that they are internal in the sense of Proposition 2.6 of (leonard2014survey).
@@
A.6.7 Proposition 3.6
Proof of Proposition 3.6.
The argument is the same as in the two-marginal case (shi2023diffusionschrodingerbridgematching, Prop. 2), except that all computations must now be performed interval by interval along the grid $t_{0}<\dots<t_{K}$. Under Assumptions A1–A3, the Doob–$h$ transform is well-defined on each interval $t_{i},t_{i+1}]$ and Lemma 11 of shi2023diffusionschrodingerbridgematching applies verbatim. The only change is that the terminal conditioning in the backward equation is at $t_{i+1}$ instead of $T$. This yields the drift
$v_{t}^{\Pi}(x)=\sigma_{t}^{2}\,\mathbb{E}_{\Pi}\!\left[\nabla\log Q^{\,t_{i},t_{i+1}}_{t_{i+1}\mid t}(X_{t_{i+1}}\mid X_{t})\,\big|\,X_{t_{i}},X_{t}\right],\qquad t\in[t_{i},t_{i+1}].$
Hence the dynamics of $\Pi$ is piecewise independent: its increment on $t_{i},t_{i+1}]$ depends only on the local bridge $Q^{t_{i},t_{i+1}}$.
The same interval-wise independence holds for any Markov $M\in\mathcal{M}$, whose SDE also factorizes on the grid. Thus both $\Pi$ and $M$ have product decompositions over the intervals, and their Radon–Nikodym derivative factorizes multiplicatively,
$\frac{d\Pi}{dM}=\prod_{i=0}^{K-1}\frac{d\Pi^{(i)}}{dM^{(i)}}.$
Taking logarithms and integrating with respect to $\Pi$ gives the additivity of the relative entropy,
$KL(\Pi\|M)=\sum_{i=0}^{K-1}KL(\Pi^{(i)}\|M^{(i)}).$
For each interval $t_{i},t_{i+1}]$, using the conditional expectation identity as in the proof of shi2023diffusionschrodingerbridgematching, we have for every $t\in[t_{i},t_{i+1}]$,
$\mathbb{E}_{\Pi_{t_{i},t}}\!\left[\big\|\sigma_{t}^{2}\,\mathbb{E}_{\Pi_{t_{i+1}\mid t_{i},t}}[\nabla\log Q^{\,t_{i},t_{i+1}}_{t_{i+1}\mid t}(X_{t_{i+1}}\mid X_{t})\mid X_{t},X_{t_{i}}]-v_{t}(X_{t})\big\|^{2}\right]$
$\geq\mathbb{E}_{\Pi_{t_{i},t}}\!\left[\big\|\sigma_{t}^{2}\mathbb{E}_{\Pi_{t_{i+1}\mid t}}[\nabla\log Q^{\,t_{i},t_{i+1}}_{t_{i+1}\mid t}(X_{t_{i+1}}\mid X_{t})\mid X_{t},X_{t_{i}}]-v_{t}^{\star}(X_{t})\big\|^{2}\right],$
where the optimal drift is defined by the orthogonal projection
$v_{t}^{\star}(x)=\sigma_{t}^{2}\,\mathbb{E}_{\Pi_{t_{i+1}\mid t}}\!\left[\nabla\log Q^{\,t_{i},t_{i+1}}_{t_{i+1}\mid t}(X_{t_{i+1}}\mid X_{t})\mid X_{t}=x_{t}\right].$
Using leonard2012, Theorem 2.3 on each interval and summing the contributions gives
$KL(\Pi\|M^{\star})=\frac{1}{2}\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\Pi_{t}}\!\left[\|v_{t}^{\Pi}(X_{t})-v_{t}^{\star}(X_{t})\|^{2}/\sigma_{t}^{2}\right]dt.$
Finally, the same Fokker–Planck uniqueness argument as in shi2023diffusionschrodingerbridgematching ensures that $M^{\star}_{t}=\Pi_{t}$ for all $t\in[t_{i},t_{i+1}]$ and all $i$. Since the grid points are included, this implies $M^{\star}=\Pi$, which concludes the proof. ∎
A.6.8 Lemma 3.1
@@
For the Markovian part, the equality follows analogously to the proof of (shi2023diffusionschrodingerbridgematching).
For each interval $t_{i},t_{i+1}]$, the same quadratic expansion gives
$2\,KL\!\left(\Pi^{(i)}\,\middle\|\,M^{(i)}\right)=2\,KL\!\left(\Pi^{(i)}\,\middle\|\,\operatorname{proj}_{\mathcal{M}}(\Pi)^{(i)}\right)+2\,KL\!\left(\operatorname{proj}_{\mathcal{M}}(\Pi)^{(i)}\,\middle\|\,M^{(i)}\right).$
Summing this identity over $i=0,\dots,K-1$, using the interval-wise independence, yields
$2\,KL(\Pi\|M)=2\,KL\!\left(\Pi\,\middle\|\,\operatorname{proj}_{\mathcal{M}}(\Pi)\right)+2\,KL\!\left(\operatorname{proj}_{\mathcal{M}}(\Pi)\,\middle\|\,M\right),$
which is the desired result.
@@
By integrating w.r.t. $\mathbb{P}$ and applying Csiszár’s Pythagorean identity (Csiszar1975, Eq. 2.6), we obtain
@@
A.6.10 Proposition 3.8
Proof of Proposition 3.8, first claim.
As a reminder, we follow the same argument as in (shi2023diffusionschrodingerbridgematching) and (debortoli2021diffusionschrodingerbridgeapplications). Applying Lemma 3.1, for any $N\in\mathbb{N}$ we obtain
@@
and boundedness $KL(\mathbb{P}^{n}\,\|\,\mathbb{P}^{\star})\leq KL(\mathbb{P}^{0}\,\|\,\mathbb{P}^{\star})<\infty.$ This proves the first claim. ∎
Proof of Proposition 3.8, second claim.
We proceed by induction, adapting the argument of (debortoli2021diffusionschrodingerbridgeapplications, Appendix C.8).
@@
and $P^{\star}$ is indeed the weak solution produced by the IMFF algorithm, proving the second claim. ∎
A.6.11 Theorem 3.2
Proof of Theorem 3.2.
As a reminder, the argument is the same as in (shi2023diffusionschrodingerbridgematching) and (debortoli2021diffusionschrodingerbridgeapplications), but adapted to the multi-marginal setting.
By lower semicontinuity of KL, this implies
$\lim_{n\to\infty}KL(\mathbb{P}^{n}\,\|\,\mathbb{P}^{\star})=0.$
Finally, the inequality